# Chiral excitation flows of multinode network based on synthetic gauge fields

### v1

@@ Abstract
Chiral excitation flows have drawn a lot of attention for their unique unidirectionality. Such flows have been studied in three-node networks with synthetic gauge fields (SGFs), while they are barely realized as the number of nodes increases. In this work, we propose a scheme to achieve chiral flows in $n$-node networks, where an auxiliary node is introduced to govern the system. This auxiliary node is coupled to all the network nodes, forming sub-triangle structures with interference paths in these networks. We find the implicit chiral symmetry behind the perfect chiral flow and propose the universal criteria that incorporate previous models, facilitating the implementation of chiral transmission in various networks. By investigating the symmetries within these models, we present different features of the chiral flow in bosonic and spin networks. Furthermore, we extend the four-node model into a ladder network, which is promising for remote state transfer in practical systems with less complexity. Our scheme can be realized in state-of-the-art experimental systems, such as superconducting circuits and magnetic photonic lattices, thereby opening up new possibilities for future quantum networks.
@@ I Introduction
The chiral excitation flow is a distinguished form of population transfer observed in quantum systems, exhibiting a unidirectional circulation characterized by chirality. A notable example of continuous chiral flows is dissipation-free edge currents in the quantum Hall effect [1, 2, 3], which was first observed under a perpendicular magnetic field [4, 5] and later related to non-trivial topological effects [6]. With the advancement of quantum technologies, edge currents can also be generated with synthetic gauge fields [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17] in photonic systems [18, 19] or other artificial systems [20, 21]. On the other hand, a discrete version of chiral excitation flows can also occur in artificial few-body systems characterized by broken time-reversal symmetry, parity symmetry, or both [22, 23, 24]. In addition, they can serve as fundamental building blocks for creating larger artificial lattice models to study quantum Hall effects, such as the integer [18] and fractional quantum Hall states [22].
@@
The quantum network connects various quantum systems via proven-secure channels and supports quantum state transfer between separated nodes, which is potentially one of the most wide-ranging applications of current quantum technologies [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37]. In practice, the nodes designed for storing and processing quantum information in quantum networks can be realized with superconducting qubits [38, 39, 40, 41, 42], atomic systems [43, 26, 44, 45], quantum dots [46, 47], or spin vacancies [48, 49, 50, 51]. Between these nodes, quantum channels play the role of bridges to coherently transfer information. In this framework, a chiral flow can provide site-by-site chiral transmission and enable unidirectional transfer of quantum states or entanglement in large-scale quantum networks. The chiral flow in a few-node network has been theoretically studied in various quantum systems, such as superconducting circuits, photonic lattices [52], Rydberg-atom arrays [53], giant atoms [54], and cavity-magnonic systems [55], and even experimentally achieved in superconducting circuits [22, 23, 24]. However, these works demonstrate perfect chiral flows only in a triangular three-node network and fail to work in larger networks. Whether perfect chiral excitation flows exist in multinode networks is still an open question, which is crucial for further applications in lattices with non-triangular geometries.
@@
In this work, we develop a scheme to generate perfect chiral excitation flows within $n$-node networks based on synthetic gauge fields (SGFs), where we introduce an auxiliary node and propose a new mode for realizing the chiral dynamics. Beyond previous studies, we find the universal criteria for conducting chiral flows in different systems. For a three-node boson-SGF network, the broken time-reversal symmetry plays a critical role in the system’s chiral dynamics, leading to complex hoppings between nodes, where the hopping phases cannot be simultaneously eliminated by a gauge transformation [10]. By introducing an auxiliary node, we study the node-assisted SGF (aSGF) model with broken time-reversal (TR) symmetry in each sub-structure of the system, in which the inherent relation between the $\pi/2$ hopping phase and a perfect chiral flow is unveiled. More generally, our aSGF model enables the emergence of these chiral flows within the $n$-node network by appropriately designing additional hoppings. In contrast to the boson system, when we generalize the SGF model to the spin or two-level systems, where TR symmetry is preserved, chiral flows can still occur but exhibit opposite chiral dynamics for up and down states. As a result, we study the fundamental elements for different models that support perfect chiral flows in the single or multi-excitation subspace and propose the universal criteria that incorporate previous models: (i) the symmetric and equally spaced energy spectrum in the corresponding excitation subspace and (ii) a complete set of chiral modes. In addition, we build an extended four-node network to function as a larger scale network that also features chirality and can potentially be employed for directional remote transmission. To make these models feasible in practice, we investigate the influences of different types of imperfections in the aSGF network and introduce two experimental platforms for implementing our model, including superconducting circuits and magnetic photonic lattice, which contribute to the chiral state transfer and directional entanglement distribution for future quantum network and quantum simulation applications.
@@
The remainder of the paper is organized as follows. First, in Sec. II, we introduce the boson-SGF model and show the collapse of chiral excitation flow when generalizing previous studies to the four-node case. Thus, in Sec. III, we propose the aSGF model to rebuild the chiral flow in the $n$-node network by introducing an auxiliary node. We then investigate the spin-SGF model that preserves TR symmetry and supports opposite chiral dynamics, and derive the general criteria for different models to conduct chiral flows. In Sec. IV, we horizontally expand the four-node network to a large ladder, establishing a chiral flow between remote nodes with less complexity. Furthermore, in Sec. V, we discuss the sensitivity of the chiral flow to the imperfections in a four-node boson-SGF network, where the parameters are taken from experimental data. Also, the experimental feasibility of our scheme is proposed in Sec. VI. Finally, in Sec. VII, we summarize the main conclusion of this work and outlook for future applications.
@@ II.1 Synthetic gauge field model
To illustrate the concept of the chiral excitation flow, we first discuss the three-node bosonic network, as shown in Fig. 1(a). In this configuration, three boson modes with the identical frequency $\omega_{0}$ are positioned at the three nodes and coupled with their nearest-neighbor (NN) sites, forming a triangle arrangement in the real space. Each mode can be described by the annihilation/creation operators $a_{j},a^{\dagger}_{j}$, where $j=1,2,3$ labels each node site, and the neighboring modes are coupled via the complex hoppings $J_{jk}=J_{0}e^{i\theta_{jk}}$. In the interaction picture, this three-node network is described by the following tight-binding Hamiltonian ($\hbar=1$)
@@
where the hopping phases $\theta_{jk}$ are induced from the SGF [22, 53]. We refer to this model as a three-node boson-SGF network, which can be easily extended to the $n$-node case where the nodes are arranged in a regular polygon.
Figure 1: (a) Illustration of the three-node boson-SGF model, where the synthetic flux $\theta_{3}$ determines the dynamical chirality. The evolution of excited-state populations $P_{e}^{j}(t)$ is shown for (b) $\theta_{3}=\pi$ and (c) $\theta_{3}=\pi/2$. The chiral excitation flow emerges when the synthetic flux of the loop $1\rightarrow 2\rightarrow 3$ is half-odd-integer multiples of $\pi$, yet the chirality disappears when the system preserves TR symmetry, i.e., $\theta_{3}=\mathbb{Z}\pi$. (d) Illustration of the four-node boson-SGF model. Similarly, (e) and (f) show the populations for $\theta_{4}=\pi$ and $\theta_{4}=\pi/2$, respectively. There is no chirality in the dynamics regardless of whether the synthetic flux $\theta_{4}$ breaks TR symmetry or not.
@@ II.2 Chiral flow in the single-excitation subspace
The three-node chiral excitation flow has been investigated in different models [22, 23, 24, 52]. Here, we focus on the single-excitation subspace of the boson-SGF model, where the 1-st node is initially prepared in its excited state $|1\rangle$ and others are in their ground state $|0\rangle$. The wavefunction of the system is written as $|\psi(t)\rangle=\sum_{j=1}^{n}C_{e}^{j}(t)a^{\dagger}_{j}|G\rangle$, with initial condition $C_{e}^{j}(0)=\delta_{j1}$, where $|G\rangle\equiv\otimes_{j=1}^{n}|0_{j}\rangle$ is the ground state of the system. From the Hamiltonian in Eq. (1), we numerically calculate and plot the time evolution of excited-state populations of the three bosonic modes. As shown in Fig 1(b), when the synthetic flux through the three-node loop is an integer multiple of $\pi$, i.e., $\theta_{3}=\mathbb{Z}\pi$, the system has TR symmetry [10], and the excitation will flow in both directions (clockwise and anti-clockwise) with equal probability. However, when TR symmetry is broken, the excitation flow tends to perform chirality. Particularly, Fig. 1(c) shows that when $\theta_{3}$ is a half-integer multiples of $\pi$, i.e., $\theta_{3}=(2\mathbb{Z}+1)\pi/2$, a perfect chiral excitation flow emerges. In this case, the excited-state populations $P^{j}_{e}(t)=|C_{e}^{j}(t)|^{2}$ can be analytically solved [52]:
@@
It takes the dimensionless time $J_{0}T_{0}=2\pi/3\sqrt{3}\approx 1.21$ to fully transfer the excited-state population from site $j$ to site $(j+1)~{}\text{mod}~{}3$. Furthermore, the dynamics of the excited-state populations for two nearest-neighboring nodes, e.g., $P^{1}_{e}(t)$ and $P^{2}_{e}(t)$, turns out to have a $32\pi/$ 3 phase shift. This feature explicitly interprets Fig. 1(c) and implies the dynamical chirality of this network. We also note that as $\theta_{3}=(2\mathbb{Z}+1)\pi/2$, the system exhibits a symmetric spectrum in the single-excitation subspace with energy levels $E=\{-\sqrt{3}J_{0},0,\sqrt{3}J_{0}\}$, which reveals an underlying chiral symmetry. We will revisit this point frequently when explaining and designing multinode chiral flows in Sec. III.
Nevertheless, the SGF model fails to achieve such chiral flows in networks consisting of four or more nodes. In Figs. 1(e) and 1(f), we plot the time evolution of the populations for a four-node SGF network. When the synthetic flux through this four-node loop ($\theta_{4}=\pi/2$) breaks TR symmetry, the excitation dynamics still shows no chirality. In fact, the populations of the 2-nd node and 4-th node are always the same, $P^{2}_{e}(t)=P^{4}_{e}(t)$, with details and analytical proof given in Appendix A. If the synthetic flux $\theta_{4}=\pi$, the population of the 3-rd node vanishes, and thus it can be referred to as a “dark site” [56, 57]. This results from the Aharonov-Bohm effect due to the destructive quantum interference from path $1\rightarrow 2\rightarrow 3$ and path $1\rightarrow 4\rightarrow 3$. Further, as discussed in Appendix B, theoretical analysis suggests that the general $n$-node ($n\geq 4$) SGF model cannot support a perfect chiral flow.
@@ III Chiral flows in multinode networks
III.1 Perfect chiral flow assisted with auxiliary node in a four-node network
Motivated by the quantum interference of the two paths in the four-node SGF model, we here introduce an auxiliary node that is positioned at the center of the network and couples to all network nodes, dividing the network into four sub-triangles, as depicted in Fig. 2(a). In this case, with appropriate parameters, we demonstrate that the quantum interference assisted by sub-triangles from different paths is able to induce a perfect chiral excitation flow.
The central-located auxiliary node can be described by annihilation and creation operators $c$ and $c^{\dagger}$. For simplicity, this auxiliary node is assumed to homogeneously couple to other nodes with a real-valued hopping rate. The Hamiltonian of the node-assisted SGF (aSGF) model reads
@@ III.1 Perfect chiral flow assisted with auxiliary node in a four-node network
where $\beta_{c}$ characterizes the hopping strength between the auxiliary node and other network nodes. This aSGF Hamiltonian will return to $H_{\rm SGF}^{4}$ as $\beta_{c}=0$. Here, we focus on the single-excitation subspace, and the system’s time-dependent wavefunction can be written as $|\psi(t)\rangle=\sum_{j=1}^{N}\left[C_{e}^{j}(t)a_{j}^{\dagger}+C_{e}^{c}(t)c^{\dagger}\right]|G\rangle$, where $|G\rangle\equiv(\otimes_{j=1}^{N}|0_{j}\rangle)\otimes|0_{c}\rangle$ is the new ground state, and we set the initial state as $C_{e}^{j}(0)=\delta_{j1}$ and $C_{e}^{c}(0)=0$. The dynamics of the system can be analytically solved, and a detailed derivation can be found in Appendix C. Remarkably, a perfect chiral flow occurs when the central coupling strength is given by $\beta_{c}=2$ as well as the synthetic flux through the four-node loop being $\theta_{4}=2\pi$. This optimal value can be directly found in the solution of excited-state populations $P_{e}^{j}=|C_{e}^{j}(t)|^{2}$, where
@@
When $\beta_{c}=2$, the two cosine functions in Eq. (4) can simultaneously reach the peak value, meaning the complete transmission of the excited-state population between neighboring nodes. Particularly, this configuration of hopping phases ($\theta_{j,j+1}=\pi/2$ and the hopping phase to the auxiliary node is 00) can still be interpreted as a vertical uniform magnetic field with a symmetric gauge chosen. Gauge invariance can be verified numerically: If one works in Landau gauge instead, $\mathbf{A}=(-By,0,0)$, the same dynamical results will be obtained.
@@
Figure 2: (a) The four-node network assisted by an auxiliary node, referred to as the four-node boson-aSGF model. Each network node is connected to the auxiliary node with the coupling $\beta_{c}J_{0}$. (b) Evolution of the excited-state populations for $\theta_{4}=2\pi$. In each sub-triangle structure, TR symmetry is broken with the synthetic flux $\theta_{\vartriangle}=\pi/2$. (c) Evolution of the excited-state populations for $\theta_{4}=\pi/2$. In contrast to the three-node SGF model, the dynamics shows no chirality with the $\pi/2$ total flux.
@@
To intuitively understand the underlying cause for the chiral flow that emerged in this auxiliary-mediated network, it is instructive to focus on sub-triangles inside the four-node pattern. The pure four-node boson-SGF model in Fig. 1(d) does not support a chiral flow due to the absence of the triangular sub-structures. However, as shown in Fig. 2(a), the auxiliary node can provide four sub-triangles that share the same hopping phase (or flux) $\theta_{\vartriangle}=\theta_{4}/4$. When we set $\theta_{\vartriangle}$ as the phase that totally breaks TR symmetry in the three-node boson-SGF model, e.g., $\theta_{\vartriangle}=\pi/2$, the perfect chiral flow between these four nodes indeed occurs, which implies that the three-node structure acts as a fundamental unit dominating the expected four-node chiral flow. In contrast, in Fig. 2(c), we show a completely different scenario where the chiral flow collapses as $\theta_{4}=\pi/2$. This demonstrates that achieving a perfect chiral flow requires broken TR symmetry in each sub-triangle rather than in the whole network. In fact, the hopping phase $\theta_{j,j+1}=\pi/2$ ensures a chiral symmetry that facilitates the chiral flow, as discussed in the following subsection.
@@
Further, the auxiliary node also contributes to the formation of chiral flow by mediating the energy spectrum and removing degeneracy. The energy levels of the four-node boson-SGF model are given by $E_{1,\pm}=\pm 2J_{0}$ and $E_{0,\pm}=0$, which results in the imperfect oscillations at even sites (see Appendix B for details), as shown in Fig. 1(e). Due to the internal chiral symmetry of $H^{4}_{\rm SGF}$, the degenerate zero-energy eigenstates are in pair, given by $|\varphi_{0,+}\rangle=(1,0,1,0)^{T}/\sqrt{2}$, $|\varphi_{0,-}\rangle=(0,1,0,1)^{T}/\sqrt{2}$. However, in the presence of the auxiliary node, these two states are lifted, along with the emergence of a new zero-energy state. The chiral symmetry also persists as the auxiliary node is introduced, with the single-particle form of chiral symmetry operator $\mathcal{C}$ written as
@@
where $\mathcal{H}_{\mathrm{aSGF}}^{4}$ is the single-particle Hamiltonian:
@@
Thus, the non-zero energy eigenstates emerge in pair $\{|\varphi\rangle,\mathcal{C}|\varphi\rangle\}$ with opposite eigenenergies, while the sole unpaired state is the new zero-energy eigenstate [58]. In Eq. (4), the lifted energies $E_{\beta_{c},\pm}=\pm 2\beta_{c}J_{0}$ result in an extra standing wave to this network, and the newly generated zero-energy state is denoted by $|\varphi_{0}^{\prime}\rangle=\sum_{j}^{n}\sigma_{j}^{+}|G\rangle/\sqrt{n}$, contributing the constant $(-1)^{j}/4$ as described in the last term. In order to ascertain the value of $\beta_{c}$ and validate the feasibility of conducting the chiral flow from a physical perspective, we rewrite the probability amplitudes in Eq. (4) as

### v2

# Chiral excitation flows of a multinode network based on synthetic gauge fields

@@ Abstract
Chiral excitation flows have attracted significant attention due to their unique unidirectionality. Such flows have been studied in three-node networks with synthetic gauge fields (SGFs), but the general theory of chiral flows in multinode networks requires further research and development. In this work, we propose a scheme to achieve chiral flows in $n$-node networks, where an auxiliary node is introduced to govern the system. This auxiliary node is coupled to all the network nodes, forming subtriangle structures with interference paths in these networks. We find the implicit chiral symmetry behind the perfect chiral flow and propose the universal criteria that incorporate previous models, facilitating the implementation of chiral transmission in various networks. By investigating the symmetries within these models, we present different features of chiral flows in bosonic and spin networks. Furthermore, we extend the four-node model into a ladder network, which is promising for remote state transfer in practical systems with reduced complexity. Our scheme can be realized in state-of-the-art experimental systems, such as superconducting circuits, magnetic photonic lattices, and ultracold atoms, thereby opening up possibilities for future quantum networks.
@@ I Introduction
Chiral excitation flow is a distinctive form of population transfer observed in quantum systems, that exhibits a unidirectional circulation characterized by chirality. A notable example of continuous chiral flows is dissipation-free edge currents in the quantum Hall effect [1, 2, 3], which was first observed under a perpendicular magnetic field [4, 5] and later related to nontrivial topological effects [6]. With the advancement of quantum technologies, edge currents can also be generated with synthetic gauge fields [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17] in photonic systems [18, 19, 20], ultracold quantum gases [21, 22, 23, 24, 25, 26] or other artificial systems [27, 28]. A discrete version of chiral excitation flows can also occur in artificial few-body systems characterized by broken time-reversal symmetry, parity symmetry, or both [29, 30, 31]. In addition, they can serve as fundamental building blocks for creating larger artificial lattice models to study quantum Hall effects, such as the integer [18] and fractional quantum Hall states [29].
@@
Quantum networks connect various quantum systems via proven-secure channels and support quantum state transfer between separated nodes, which is potentially one of the most wide-ranging applications of current quantum technologies [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44]. In practice, the nodes designed for storing and processing quantum information in quantum networks can be realized with superconducting qubits [45, 46, 47, 48, 49], atomic systems [50, 33, 51, 52], quantum dots [53, 54], or spin vacancies [55, 56, 57, 58]. Between these nodes, quantum channels play the role of bridges to coherently transfer information. In this framework, a chiral flow can provide site-by-site chiral transmission and enable unidirectional transfer of quantum states or entanglement in large-scale quantum networks. Chiral flows in few-node networks have been theoretically studied in various quantum systems, including superconducting circuits, photonic lattices [59], Rydberg-atom arrays [60], giant atoms [61], and cavity-magnonic systems [62], and they have even been experimentally achieved in both superconducting circuits [29, 30, 31] and trapped-ion quantum processor[63]; however, these works demonstrated perfect chiral flows only in a triangular three-node network and fail to operate in larger networks. The systematic implementation of perfect chiral excitation flows in multinode networks—which could potentially lead to a wide range of applications in networks beyond a single triangular geometry—remains largely unexplored.
@@
In this work, we develop a scheme to generate perfect chiral excitation flows within $n$-node networks based on synthetic gauge fields (SGFs), introducing an auxiliary node and proposing a mode for realizing the chiral dynamics. Moving beyond previous studies, we find the universal criteria for conducting chiral flows in different systems. For a three-node boson-SGF network, the broken time-reversal (TR) symmetry plays a critical role in the system’s chiral dynamics, leading to complex hoppings between nodes, where the hopping phases cannot be simultaneously eliminated by a gauge transformation [10]. By introducing an auxiliary node, we study the node-assisted SGF (aSGF) model with broken TR symmetry in each substructure of the system, in which the inherent relation between the $\pi/2$ hopping phase and a perfect chiral flow is unveiled. More generally, our aSGF model enables the emergence of these chiral flows within the $n$-node network by appropriately designing additional hoppings. In contrast to the bosonic system, when we generalize the SGF model to spin or two-level systems, where TR symmetry is preserved, chiral flows can still occur but exhibit opposite chiral dynamics for the up and down states. As a result, we study the fundamental elements for different models that support perfect chiral flows in the single or multiexcitation subspace and propose universal criteria that incorporate previous models: (i) the symmetric and equally spaced energy spectrum in the corresponding excitation subspace and (ii) a complete set of chiral modes. In addition, we build an extended four-node network to function as a larger-scale network that also features chirality and can potentially be employed for directional remote transmission. To make these models feasible in practice, we investigate the influences of different types of imperfections in the aSGF network and introduce three experimental platforms for implementing our model: superconducting circuits, magnetic photonic lattices , and ultracold atoms. These platforms contribute to the chiral state transfer and directional entanglement distribution for future quantum network and quantum simulation applications.
@@
The remainder of theis paper is organized as follows. First, in Sec. II, we introduce the boson-SGF model and show the collapse of chiral excitation flow when generalizing previous studies to the four-node case. Thus, in Sec. III, we propose the aSGF model to rebuild the chiral flow in an $n$-node network by introducing an auxiliary node. We then investigate the spin-SGF model that preserves TR symmetry and supports opposite chiral dynamics and derive the general criteria for different models to conduct chiral flows. In Sec. IV, we horizontally expand the four-node network to a large ladder, establishing a chiral flow between remote nodes with reduced complexity. In Sec. V, we discuss the sensitivity of the chiral flow to the imperfections in a four-node boson-SGF network with parameters taken from experimental data. The experimental feasibility of our scheme is considered in Sec. VI. Finally, in Sec. VII, we summarize the main conclusions of this work and present the outlook for future applications.
@@ II.1 Synthetic gauge field model
To illustrate the concept of the chiral excitation flow, we first discuss the three-node bosonic network, as shown in Fig. 1(a). In this configuration, three boson modes with the identical frequency $\omega_{0}$ are positioned at the three nodes and coupled with their nearest-neighbor (NN) sites, forming a triangle arrangement in real space. Each mode can be described by the annihilation and creation operators $a_{j}$ and $a^{\dagger}_{j}$, where $j=1,2,3$ labels each node site and the neighboring modes are coupled via the complex hoppings $J_{jk}=J_{0}e^{i\theta_{jk}}$. In the interaction picture, this three-node network is described by the following tight-binding Hamiltonian ($\hbar=1$)
@@
where the hopping phases $\theta_{jk}$ are induced from the SGF [29, 60]. We refer to this model as a three-node boson-SGF network, which can be readily extended to the $n$-node case with nodes arranged in a regular polygon [e.g., see Figs. 1(d), 10(a), and 10(b) for $n=4,5$, and $6$, respectively].
Figure 1: (a) Illustration of the three-node boson-SGF model, where the synthetic flux $\theta_{3}$ determines the dynamical chirality. The evolution of excited-state populations $P_{e}^{j}(t)$ is shown for (b) $\theta_{3}=\pi$ and (c) $\theta_{3}=\pi/2$. The chiral excitation flow emerges when the synthetic flux of the loop $1\rightarrow 2\rightarrow 3$ is half-odd-integer multiples of $\pi$, yet the chirality disappears when the system preserves TR symmetry, i.e., $\theta_{3}=p\pi$ ($p\in\mathbb{Z}$). (d) Illustration of the four-node boson-SGF model. Similarly, (e) and (f) show the populations for $\theta_{4}=\pi$ and $\theta_{4}=\pi/2$, respectively. There is no chirality in the dynamics regardless of whether the synthetic flux $\theta_{4}$ breaks TR symmetry.
@@ II.2 Chiral flow in the single-excitation subspace
The three-node chiral excitation flow has been investigated in different models [29, 30, 31, 59]. Here, we focus on the single-excitation subspace of the boson-SGF model, where the first node is initially prepared in its excited state $|1\rangle$ and others are in their ground state $|0\rangle$. The wave function of the system is written as $|\psi(t)\rangle=\sum_{j=1}^{n}C_{e}^{j}(t)a^{\dagger}_{j}|G\rangle$, with initial condition $C_{e}^{j}(0)=\delta_{j1}$, where $|G\rangle\equiv\otimes_{j=1}^{n}|0_{j}\rangle$ is the ground state of the system. From the Hamiltonian in Eq. (1), we numerically calculate and plot the time evolution of the excited-state populations of the three bosonic modes. As shown in Fig 1(b), when the synthetic flux through the three-node loop is an integer multiple of $\pi$, i.e., $\theta_{3}=p\pi$ ($p\in\mathbb{Z}$), the system has TR symmetry [10], and the excitation will flow in both directions (clockwise and anticlockwise) with equal probability; however, when TR symmetry is broken, the excitation flow tends to be chiral. In particular, Fig. 1(c) shows that when $\theta_{3}$ is a half-integer multiples of $\pi$, i.e., $\theta_{3}=(p+1/2)\pi$ with $p\in\mathbb{Z}$, a perfect chiral excitation flow emerges. In this case, the excited-state populations $P^{j}_{e}(t)=|C_{e}^{j}(t)|^{2}$ can be analytically solved [59, 64]:
@@
It takes the dimensionless time $J_{0}T_{0}=2\pi/3\sqrt{3}\approx 1.21$ to fully transfer the excited-state population from site $j$ to site $(j+1)~{}\text{mod}3$. Furthermore, the dynamics of the excited-state populations for two nearest-neighboring nodes, e.g., $P^{1}_{e}(t)$ and $P^{2}_{e}(t)$, turns out to have a $32\pi/$ 3 phase shift. This feature explicitly interprets Fig. 1(c) and implies the dynamical chirality of this network. We also note that as $\theta_{3}=(p+1/2)\pi$ in which $p\in\mathbb{Z}$, the system exhibits a symmetric spectrum in the single-excitation subspace with energy levels $E=\{-\sqrt{3}J_{0},0,\sqrt{3}J_{0}\}$, which reveals an underlying chiral symmetry, namely, $\mathcal{C}^{-1}\mathcal{H}\mathcal{C}=-\mathcal{H}$. This usage of the term “chiral” differs from its previous context in “chiral flow”, where “chiral” refers to the lack of inversion symmetry in excitation flows. We will revisit this point frequently when explaining and designing multinode chiral flows in Sec. III.
However, after a simple generalization, the SGF model fails to achieve such chiral flows in networks consisting of four or more nodes [64, 30]. In Figs. 1(e) and 1(f), we plot the time evolution of the populations for a four-node SGF network. When the synthetic flux through this four-node loop ($\theta_{4}=\pi/2$) breaks TR symmetry, the excitation dynamics still show no chirality. In fact, the populations of the second and fourth nodes are always the same, $P^{2}_{e}(t)=P^{4}_{e}(t)$, with details and analytical proof given in Appendix A. If the synthetic flux $\theta_{4}=\pi$, the population of the third node vanishes, and it can thus be referred to as a “dark site” [65, 66]. This results from the Aharonov-Bohm effect due to the destructive quantum interference from the paths $1\rightarrow 2\rightarrow 3$ and $1\rightarrow 4\rightarrow 3$. Further, as discussed in Appendix B, theoretical analysis suggests that the general $n$-node ($n\geq 4$) SGF model cannot support a perfect chiral flow.
@@ III Chiral flows in multinode networks
III.1 Perfect chiral flow assisted by an auxiliary node in a four-node network
Motivated by the quantum interference of the two paths in the four-node SGF model, here, we introduce an auxiliary node that is positioned at the center of the network and couples to all network nodes, dividing the network into four subtriangles, as depicted in Fig. 2(a). In this case, with appropriate parameters, we demonstrate that the quantum interference assisted by subtriangles from different paths is able to induce a perfect chiral excitation flow.
The centrally located auxiliary node can be described by annihilation and creation operators $c$ and $c^{\dagger}$. For simplicity, this auxiliary node is assumed to homogeneously couple to other nodes with a real-valued hopping rate. The Hamiltonian of the node-assisted SGF (aSGF) model reads
@@ III.1 Perfect chiral flow assisted by an auxiliary node in a four-node network
where $\beta_{c}$ characterizes the hopping strength between the auxiliary node and other network nodes. This aSGF Hamiltonian will return to $H_{\rm SGF}^{4}$ as $\beta_{c}=0$. Here, we focus on the single-excitation subspace, and the system’s time-dependent wave function can be written as $|\psi(t)\rangle=\sum_{j=1}^{N}\left[C_{e}^{j}(t)a_{j}^{\dagger}+C_{e}^{c}(t)c^{\dagger}\right]|G\rangle$, where $|G\rangle\equiv(\otimes_{j=1}^{N}|0_{j}\rangle)\otimes|0_{c}\rangle$ is the new ground state, and we set the initial state as $C_{e}^{j}(0)=\delta_{j1}$ and $C_{e}^{c}(0)=0$. The dynamics of the system can be analytically solved, and a detailed derivation can be found in Appendix C. Remarkably, a perfect chiral flow occurs when the central coupling strength is given by $\beta_{c}=2$ and the synthetic flux through the four-node loop is $\theta_{4}=2\pi$. This optimal value can be directly found in the solution of excited-state populations $P_{e}^{j}=|C_{e}^{j}(t)|^{2}$, where
@@
When $\beta_{c}=2$, the two cosine functions in Eq. (4) can simultaneously reach their peak values, indicating the complete transmission of the excited-state population between neighboring nodes. In particular, this configuration of hopping phases ($\theta_{j,j+1}=\pi/2$, and the hopping phase to the auxiliary node is 00) can still be interpreted as a vertical uniform magnetic field with a symmetric gauge chosen. Gauge invariance can be verified numerically: If one instead works in Landau gauge, $\bm{A}=(-By,0,0)$, the same dynamical results will be obtained.
@@
Figure 2: (a) Four-node network assisted by an auxiliary node, referred to as the four-node boson-aSGF model. Each network node is connected to the auxiliary node with the coupling $\beta_{c}J_{0}$. (b) Evolution of the excited-state populations for $\theta_{4}=2\pi$. In each subtriangle structure, TR symmetry is broken with the synthetic flux $\theta_{\vartriangle}=\pi/2$. (c) Evolution of the excited-state populations for $\theta_{4}=\pi/2$. In contrast to the three-node SGF model, the dynamics show no chirality with the $\pi/2$ total flux.
@@
To intuitively understand the underlying cause of the chiral flow that emerges in this auxiliary-mediated network, it is instructive to focus on subtriangles inside the four-node pattern. The pure four-node boson-SGF model in Fig. 1(d) does not support a chiral flow due to the absence of triangular substructures; however, as shown in Fig. 2(a), the auxiliary node can provide four subtriangles that share the same hopping phase (or flux) $\theta_{\vartriangle}=\theta_{4}/4$. When we set $\theta_{\vartriangle}$ as the phase that completely breaks TR symmetry in the three-node boson-SGF model, e.g., $\theta_{\vartriangle}=\pi/2$, the perfect chiral flow between these four nodes indeed occurs, which implies that the three-node structure acts as a fundamental unit dominating the expected four-node chiral flow. In contrast, in Fig. 2(c), we show a completely different scenario in which the chiral flow collapses as $\theta_{4}=\pi/2$. This demonstrates that achieving a perfect chiral flow requires broken TR symmetry in each subtriangle rather than in the whole network. In fact, the hopping phase $\theta_{j,j+1}=\pi/2$ ensures a chiral symmetry that facilitates the chiral flow, as discussed in the next subsection.
@@
Further, the auxiliary node also contributes to the formation of chiral flow by mediating the energy spectrum and removing degeneracy. The energy levels of the four-node boson-SGF model are given by $E_{1,\pm}=\pm 2J_{0}$ and $E_{0,\pm}=0$, which results in imperfect oscillations at even sites (see Appendix B for details), as shown in Fig. 1(e). Due to the internal chiral symmetry of $H^{4}_{\rm SGF}$, the degenerate zero-energy eigenstates are in pairs, given by $|\varphi_{0,+}\rangle=(1,0,1,0)^{T}/\sqrt{2}$ and $|\varphi_{0,-}\rangle=(0,1,0,1)^{T}/\sqrt{2}$; however, in the presence of the auxiliary node, these two states are lifted, along with the emergence of a new zero-energy state. The chiral symmetry also persists as the auxiliary node is introduced, with the single-particle form of the chiral symmetry operator $\mathcal{C}$ written as
@@
where $\mathcal{H}_{\mathrm{aSGF}}^{4}$ represents the single-particle Hamiltonian [see Eq. (45) with $\theta_{j,j+1}=\pi/2$ for the matrix form]:
@@
This configuration of hopping phases ensures the chiral symmetry of the aSGF model. The chiral transformation is equivalent to inverting the network about the diagonal “1a3” [see Fig. 2(a)] and performing the transformation $c^{\dagger}\rightarrow-c^{\dagger}$ at the same time. Under this transformation, all the hoppings reverse their directions, leading to an overall minus sign to the single-particle Hamiltonian. Thus, the nonzero energy eigenstates emerge in pairs $\{|\varphi\rangle,\mathcal{C}|\varphi\rangle\}$ with opposite eigenenergies, while the sole unpaired state should be a zero-energy eigenstate [67]. In Eq. (4), the lifted energies $E_{\beta_{c},\pm}=\pm 2\beta_{c}J_{0}$ result in an extra standing wave to this network, and the zero-energy state is denoted by $|\varphi_{0}^{\prime}\rangle=\sum_{j=1}^{4}\frac{1}{2\sqrt{2}}a_{j}^{\dagger}|G\rangle+\frac{1}{\sqrt{2}}c^{\dagger}|G\rangle$, contributing the constant $(-1)^{j-1}/4$ as described in the last term. To ascertain the value of $\beta_{c}$ and validate the feasibility of conducting the chiral flow from a physical perspective, we rewrite the probability amplitudes in Eq. (4) as